OpenStudy (anonymous):

A car with a mass of 800kg experiences a friction force of 180N and travels at a constant speed of 54km/h. Calculate the power required from the engine to travel at the constant speed.

OpenStudy (anonymous):

power=force x distance/ time.

OpenStudy (anonymous):

does the friction force stay the same at any velocity?

OpenStudy (apoorvk):

If it's moving, that means it is being acted upon by kinetic friction, and kinetic friction is constant for a moving body under the same conditions, irrespective of velocity.

OpenStudy (anonymous):

force=mass x acceleration. so how do you get the acceleration?

OpenStudy (anonymous):

force-friction force = unbalanced force and unbalanced force= mass x acceleration. the acceleration is 0 because he is travelling at a constant speed. so f - 180= 800x0 then the force is 180N. am i right?

OpenStudy (anonymous):

because the car only needs to overcome the friction force to travel at the constant speed?????

OpenStudy (anonymous):

You are correct. The force provided by the engine is equal to the friction force. Recall that\[Power = {Force \times Distance \over Time}\]but\[{Distance \over Time} = Velocity\]therefore\[Power = Force \times Velocity\]